OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−4, −1) and (4, 5)?

OpenStudy (anonymous):

First, find the slope of this line. THe slope is 3/4 Then write it in the form y = mx + b y = 3/4x + b Then plug in one of the points. -1 = 3/4(-4) + b b = 2. Therefore, the y-intercept is 2